# GroundingDINO-US-SAM: Text-Prompted Multi-Organ Segmentation in Ultrasound with LoRA-Tuned Vision–Language Models

### v2

@@ Abstract
Accurate and generalizable object segmentation in ultrasound imaging remains a significant challenge due to anatomical variability, diverse imaging protocols, and limited annotated data. In this study, we propose a prompt-driven vision-language model (VLM) that integrates Grounding DINO with SAM2 to enable object segmentation across multiple ultrasound organs. A total of 18 public ultrasound datasets, encompassing the breast, thyroid, liver, prostate, kidney, and paraspinal muscle, were utilized. These datasets were divided into 15 for fine-tuning and validation of Grounding DINO using Low Rank Adaptation (LoRA) to the ultrasound domain, and 3 were held out entirely for testing to evaluate performance in unseen distributions. Comprehensive experiments demonstrate that our approach outperforms state-of-the-art segmentation methods, including UniverSeg, MedSAM, MedCLIP-SAM, BiomedParse, and SAMUS on most seen datasets while maintaining strong performance on unseen datasets without additional fine-tuning. These results underscore the promise of VLMs in scalable and robust ultrasound image analysis, reducing dependence on large, organ-specific annotated datasets. We will publish our code on code.sonography.ai after acceptance.
@@ 1 Introduction
Despite the numerous advancements mentioned above, such as reducing annotation needs or enhancing robustness across different anatomical structures, these methods often address isolated challenges and fall short of fulfilling broader clinical and technical requirements. A universally generalizable solution that performs reliably across a wide range of tasks, organs, and imaging conditions remains elusive. This has fueled growing interest in more flexible and adaptable paradigms, such as prompt-driven segmentation using vision-language models (VLMs). VLMs offer several advantages: they enable zero-shot generalization to previously unseen anatomical structures through text prompts, allow seamless adaptation across different segmentation tasks without task-specific retraining, and promote scalability across imaging modalities and clinical use cases.
Prompt-driven segmentation using VLM approaches has emerged as a promising alternative for generalizing to previously unseen anatomical structures, enabling segmentation without the need for additional training [14]. Building on this paradigm, several models have been proposed to leverage the strengths of VLM integration for more generalizable medical image segmentation. For example, Grounding DINO [15] is one of the widely used VLM approaches that performs open-set object detection guided by input prompts. By combining image features with text embeddings, it enables precise localization of objects described in free-form text, making it one of the leading approaches for language-guided visual understanding. BiomedParse [16] demonstrates the feasibility of joint segmentation, detection, and recognition across multiple modalities, reinforcing the potential of foundation models. Variants such as MedCLIP-SAM [17] and MedCLIP-SAM2 [18] integrate CLIP-style [19] image-text pre-training with segmenta anything model (SAM) [20] for CT and MRI segmentation. However, their effectiveness on ultrasound data remains limited due to modality mismatches.
@@
To address the limitations of manual prompting in SAM-based segmentation, we incorporate Grounding DINO [15] as a prior model to automatically generate the bounding boxes needed for SAM-based segmentation of ultrasound images. Grounding DINO integrates cross-modal attention between image and text features to generate bounding boxes from input prompts without requiring predefined classes or user interaction. This capability enables automatic spatial prompting for SAM, allowing it to segment anatomical structures in ultrasound images based on high-level semantic cues such as organ names. Unlike prior methods like APG-SAM, which rely solely on visual detection networks and are restricted to specific anatomies (e.g., breast), our approach combines the semantic flexibility of text-based localization with the generalization ability of VLMs, resulting in a more scalable and fully automated segmentation pipeline across diverse US imaging domains. To this end, our proposed approach integrates a Low-Rank Adaptation (LoRA), tuned Grounding DINO detector, with a frozen SAM2 decoder. Notably, only 1.7% of the Grounding DINO parameters are updated during training, while the weights of SAM2 remain entirely frozen. Therefore, the proposed LoRA-based grounding strategy overcomes the geometric-only guidance of APG-SAM[26] and the rigid organ-specific prompt design of Chen et al.[28], offering a scalable, annotation-free solution for ultrasound segmentation in diverse clinical applications. The contribution of the proposed approaches can be summarized as follows:
  • Accepts free-form natural-language prompts (e.g. ‘‘tumor’’, ‘‘malignant lesion’’) without any clicks, boxes, or organ-specific retraining
@@ 2.2 Model Architecture and LoRA-Based Fine-Tuning
Figure 1: Overview of the proposed LoRA-based fine-tuning strategy integrated into the Grounding DINO framework for ultrasound image-text grounding. The architecture includes frozen backbones, LoRA-injected adapters in the feature enhancer and cross-modality decoder layers, and a SAM2 segmentation head.
As illustrated in Fig. 1, the proposed framework builds on Grounding DINO, a transformer-based vision-language model (VLM) that performs joint image-text grounding. It consists of a frozen image backbone, a frozen BERT-based text encoder, and a transformer-based encoder-decoder structure [15]. To efficiently adapt this model to ultrasound (US) data while preserving pretrained knowledge, we introduce Low-Rank Adaptation (LoRA) [49] into selected modules:
@@ 2.3 Loss Function
The proposed model is trained using a composite loss function that combines three components as outlined below. This loss supervises bounding box regression and open-vocabulary classification:
@@
1. $\mathscr{L}_{\text{L1}}$: Bounding Box Regression (Localization)
This term measures the absolute difference between predicted and ground-truth bounding box coordinates. It encourages the model to generate precise box dimensions.
2. $\mathscr{L}_{\text{GIoU}}$: Spatial Overlap (Localization)
The GIoU loss [50] extends the Intersection-over-Union (IoU) metric to penalize predictions that are far from the ground-truth boxes, even when they do not overlap. This improves convergence and spatial accuracy in object localization.
3. $\mathscr{L}_{\text{focal}}$: Contrastive Classification (Text-Guided Prediction)
Inspired by GLIP [51], we used Focal Loss [52] to supervise classification via contrastive learning between visual queries and language tokens. Specifically, the model computes dot products between each decoder query and all token embeddings from the input prompt, producing a set of logits. The focal loss is then applied to these logits to prioritize harder examples and mitigate class imbalance. This enables open-vocabulary segmentation by aligning image regions with free-form text descriptions.
This loss formulation enables the model to localize and classify objects described by text prompts with high precision while preserving generalization to unseen anatomy or phrasing.
@@ 2.5 Implementation Details
LoRA fine-tuning was conducted using a batch size of 4 and a learning rate of $5\times 10^{-4}$ with the AdamW optimizer [55] with weight decay of $1\times 10^{-4}$. Only LoRA parameters were updated, whereas all other weights remained frozen. The training loss included focal loss for classification, L1 and GIoU for localization, and a contrastive alignment loss to optimize image-text correspondence. Early stopping was employed with a patience of 20 epochs based on validation loss to mitigate overfitting. All datasets were pooled during training, and data augmentations such as random flipping, resizing, padding, erasing, and cropping were applied. This led to a single, general-purpose model that was evaluated on all datasets (seen and unseen).
@@ 3.1 Ablation Study
Table 2 summarizes the ablation study evaluating different configurations of the Grounding DINO model combined with various segmentation heads. The baseline configuration, Grounding DINO with SAM2 (trained on natural images), shows limited segmentation performance across all datasets. Fine-tuning Grounding DINO paired with MedSAM notably improves both DSC and IoU scores. However, the best performance is achieved by fine-tuning Grounding DINO paired with SAM2, which consistently outperforms the other configurations across breast, liver, and prostate datasets. Specifically, on the challenging AUL liver dataset, the DSC and IoU scores increased to $44.02\%$ and $35.79\%$, respectively. In the MicroSeg prostate dataset, fine-tuning with SAM2 yielded a DSC of $88.56\%$ and IoU of $81.17\%$, demonstrating the effectiveness of this fine-tuning strategy. On average, the fine-tuned Grounding DINO with SAM2 achieved the highest DSC of $73.99\%$ and IoU of $65.66\%$, confirming its superior generalization ability. This configuration is therefore adopted for all subsequent experiments.
Table 2: Ablation study comparing Grounding DINO’s segmentation performance (DSC and IoU scores) paired with different segmentation heads across multiple ultrasound datasets. The evaluated configurations include: (1) Grounding DINO + SAM2 (trained on natural images), (2) Fine-tuned Grounding DINO + MedSAM, and (3) Fine-tuned Grounding DINO + SAM2. Results are reported as mean $\pm$ standard deviation. Bolded results indicate the highest score in each comparison. The third method consistently outperforms other methods across all datasets and organs, and is used in the rest of the experiments.
@@
Breast BrEaST 23.87±30 18.08±27 49.62±28 37.33±24 76.65±23 66.24±23
BUID 44.09±35 36.08±35 61.43±29 49.38±25 82.92±24 75.83±25
BUSI 29.55±34 23.31±31 51.35±30 39.74±26 77.79±26 69.29±27
Liver AUL 13.77±13 7.94±8 32.50±28 23.25±23 44.02±36 35.79±33
Prostate MicroSeg 61.44±21 47.47±21 56.84±22 42.75±20 88.56±13 81.17±15
Average 34.54±26 26.58±24 50.35±27 38.49±23 73.99±24 65.66±24
@@ 3.2 Comparison on Seen Datasets
The average performance across all datasets, summarized in the last row of Table 3, shows that our method achieves the best mean DSC of $78.97\%$ and IoU of $71.13\%$, outperforming all other methods on average. This further demonstrates our approach’s strong generalization capability and effectiveness compared to existing models.
Table 3: Quantitative comparison of segmentation performance (DSC and IoU scores) on publicly available ultrasound datasets among SOTA methods: UniverSeg, BiomedParse, SAMUS, MedCLIP-SAM, MedCLIP-SAMv2, and the proposed method. Results are reported as mean $\pm$ standard deviation. Bolded results indicate the highest score in each comparison.
@@
Breast BrEaST 60.30±23 47.01±24 76.36±20 65.19±22 60.78±22 46.95±21 65.02±21 51.18±19 72.91±25 62.24±25 76.65±23 66.24±23
BUID 56.30±30 45.27±29 82.77±21 74.54±23 71.74±17 57.93±16 70.13±22 57.90±24 78.48±23 69.48±26 82.92±24 75.83±25
BUSUC 51.94±28 39.89±26 87.17±12 78.92±16 71.46±17 58.06±19 80.66±11 68.88±14 87.40±11 78.99±15 91.68±6 85.16±9
BUSUCLM 52.96±22 39.19±21 66.34±31 56.50±30 44.34±32 34.32±29 56.88±21 42.48±19 66.09±27 54.66±27 76.74±24 66.75±24
BUSB 82.58±15 72.87±20 73.93±31 66.12±31 64.40±20 50.52±20 73.84±23 62.84±24 82.99±23 75.46±23 91.04±8 84.30±11
BUSI 43.19±26 31.16±22 89.39±14 82.72±15 82.45±10 71.19±13 56.64±30 45.51±29 63.60±34 54.40±32 77.79±26 69.29±27
STU 64.73±17 50.16±19 87.07±5 77.50±8 68.46±11 53.10±12 83.77±9 73.03±13 90.11±6 82.56±10 87.89±6 78.87±9
S1 57.30±25 44.25±24 93.28±4 87.63±6 76.09±20 64.54±20 78.92±9 66.02±12 88.04±8 79.40±11 93.37±5 87.96±8
Thyroid TN3K 39.46±21 26.88±18 56.96±34 47.07±31 80.96±12 69.43±14 44.42±27 32.60±23 43.28±31 32.92±27 72.20±26 61.77±26
TG3K 37.27±19 24.57±14 1.57±5 0.88±3 82.95±6 71.22±8 67.82±17 53.50±17 70.75±20 57.73±20 80.32±9 68.04±12
@@
AUL 14.52±15 8.60±10 19.78±32 15.48±26 28.18±22 18.59±17 32.12±27 22.69±22 32.86±34 25.48±29 44.02±36 35.79±33
Prostate MicroSeg 11.62±9 6.44±6 23.98±24 16.25±20 70.84±18 57.28±18 58.67±15 42.94±14 53.04±30 41.10±25 88.56±13 81.17±15
RegPro 16.84±14 9.89±9 38.88±32 29.53±27 49.22±29 37.30±25 42.90±28 31.26±23 46.90±31 36.00±27 70.77±31 61.63±29
Kidney KidneyUS 22.22±15 13.36±10 48.03±23 34.55±20 56.37±18 41.39±18 59.86±16 44.45±15 64.93±21 51.10±20 73.49±16 60.48±19
Average 47.33±22 35.77±20 62.00±21 53.38±21 60.01±18 48.04±17 62.76±20 50.48±20 68.92±23 59.06±22 78.97±18 71.13±20

### v3

@@ Abstract
Accurate and generalizable object segmentation in ultrasound imaging remains a significant challenge due to anatomical variability, diverse imaging protocols, and limited annotated data. In this study, we propose a prompt-driven vision-language model (VLM) that integrates Grounding DINO with SAM2 (Segment Anything Model2) to enable object segmentation across multiple ultrasound organs. A total of 18 public ultrasound datasets, encompassing the breast, thyroid, liver, prostate, kidney, and paraspinal muscle, were utilized. These datasets were divided into 15 for fine-tuning and validation of Grounding DINO using Low Rank Adaptation (LoRA) to the ultrasound domain, and 3 were held out entirely for testing to evaluate performance in unseen distributions. Comprehensive experiments demonstrate that our approach outperforms state-of-the-art segmentation methods, including UniverSeg, MedSAM, MedCLIP-SAM, BiomedParse, and SAMUS on most seen datasets while maintaining strong performance on unseen datasets without additional fine-tuning. These results underscore the promise of VLMs in scalable and robust ultrasound image analysis, reducing dependence on large, organ-specific annotated datasets. We will publish our code on code.sonography.ai after acceptance.
@@ 1 Introduction
Despite the numerous advancements mentioned above, such as reducing annotation needs or enhancing robustness across different anatomical structures, these methods often address isolated challenges and fall short of fulfilling broader clinical and technical requirements. A universally generalizable solution that performs reliably across a wide range of tasks, organs, and imaging conditions remains elusive. This has fueled growing interest in more flexible and adaptable paradigms, such as prompt-driven segmentation using vision-language models (VLMs). VLMs offer several advantages: they enable zero-shot generalization to previously unseen anatomical structures through text prompts [14, 15], allow seamless adaptation across different segmentation tasks without task-specific retraining [16], and promote scalability across imaging modalities and clinical use cases [17].
Prompt-driven segmentation using VLM approaches has emerged as a promising alternative for generalizing to previously unseen anatomical structures, enabling segmentation without the need for additional training [18]. Building on this paradigm, several models have been proposed to leverage the strengths of VLM integration for more generalizable medical image segmentation 1. For example, Grounding DINO [16] is one of the widely used VLM approaches that performs open-set object detection guided by input prompts. By combining image features with text embeddings, it enables precise localization of objects described in free-form text, making it one of the leading approaches for language-guided visual understanding. BiomedParse [19] demonstrates the feasibility of joint segmentation, detection, and recognition across multiple modalities, reinforcing the potential of foundation models. Variants such as MedCLIP-SAM [15] and MedCLIP-SAM2 [20] integrate CLIP-style [14] image-text pre-training with segmenta anything model (SAM) [21] for CT and MRI segmentation. However, their effectiveness on ultrasound data remains limited due to modality mismatches.
@@
To address the limitations of manual prompting in SAM-based segmentation, we incorporate Grounding DINO [16] as a prior model to automatically generate the bounding boxes needed for SAM-based segmentation of ultrasound images. Grounding DINO integrates cross-modal attention between image and text features to generate bounding boxes from input prompts without requiring predefined classes or user interaction. This capability enables automatic spatial prompting for SAM, allowing it to segment anatomical structures in ultrasound images based on high-level semantic cues such as organ names. Unlike prior methods like APG-SAM, which rely solely on visual detection networks and are restricted to specific anatomies (e.g., breast), our approach combines the semantic flexibility of text-based localization with the generalization ability of VLMs, resulting in a more scalable and fully automated segmentation pipeline across diverse ultrasound imaging domains. To this end, our proposed approach integrates a Low-Rank Adaptation (LoRA), tuned Grounding DINO detector, with a frozen SAM2 decoder. Notably, only 1.7% of the Grounding DINO parameters are updated during training, while the weights of SAM2 remain entirely frozen. Therefore, the proposed LoRA-based grounding strategy overcomes the geometric-only guidance of APG-SAM[27] and the rigid organ-specific prompt design of Chen et al.[29], offering a scalable, annotation-free solution for ultrasound segmentation in diverse clinical applications. For natural images, Das et al. [30] demonstrated the effectiveness of prompting foundational models for omni-supervised instance segmentation, showcasing how models like Grounding DINO and SAM can be adapted to complex instance segmentation tasks. Chen et al. [31] presented a visual precision search interpretability framework built on top of Grounding DINO and SAM to study object-level model behavior. Zhang et al. [32] introduced EfficientViT‑SAM, which utilizes Grounding DINO to provide bounding box proposals to SAM in order to generate segmentation masks in a zero-shot open-world setting.
The contribution of the proposed approaches can be summarized as follows:
  • Accepts free-form natural-language prompts (e.g. "tumor", "malignant lesion") without any clicks, boxes, or organ-specific retraining
@@
Table 1: Comparison of recent prompt-driven segmentation methods on ultrasound imaging. Point and bounding box columns indicate whether the method needs manual spatial prompts. Our proposed method supports fully automated segmentation with text-based interactions, overcoming common modality mismatches and manual prompting requirements.
Model Ultrasound-based Text prompt Point Bounding box Notes
Grounding DINO Developed for natural images
BiomedParse Not optimized for US
MedCLIP-SAM Limited US performance
UltraSAM No text interaction
SAM-MedUS Requires manual spatial prompts
ClickSAM Spatial interaction needed
SAMUS Needs expert points
APG-SAM Limited to breast US
MOFO Closed organ list
Proposed Optimized for US; no manual spatial prompts needed
@@ 2.2 Model Architecture and LoRA-Based Fine-Tuning
Figure 1: The system uses frozen Grounding DINO and SAM2 for zero-shot ultrasound segmentation via natural language prompts. Image and text features are extracted, fused with attention layers, and used to predict bounding boxes. SAM2 performs segmentation based on these localized boxes.
As illustrated in Fig. 1, the proposed framework builds on Grounding DINO, a transformer-based vision-language model (VLM) that performs joint image-text grounding.
LoRA introduces low-rank trainable matrices to approximate weight updates while keeping pretrained weights frozen. Given a weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA represents the update as:
$W=W_{0}+\Delta W=W_{0}+BA,$
where $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$ are learnable matrices with rank $r\ll\min(d,k)$, significantly reducing the number of trainable parameters. $B$ is initialized to zero and $A$ is randomly initialized, so that $BA=0$ at the start of training. This ensures that the initial output of the model matches the pretrained model.
It consists of a frozen image backbone, a frozen BERT-based text encoder, and a transformer-based encoder-decoder structure [16]. To efficiently adapt this model to ultrasound data while preserving pretrained knowledge, we introduce Low-Rank Adaptation (LoRA) [53] into selected modules:
@@ 2.3 Loss Function
The proposed model is trained using a composite loss function that combines three components to jointly supervise bounding box regression, spatial alignment, and text-guided classification:
@@
The loss components are defined as follows:
1. $\mathscr{L}_{\text{L1}}$: Bounding Box Regression
This is the standard L1 loss between predicted box coordinates $\hat{b}=(x,y,w,h)$ and ground truth $b$:
$\mathscr{L}_{\text{L1}}=\lVert\hat{b}-b\rVert_{1}$ (2)
2. $\mathscr{L}_{\text{GIoU}}$: Generalized IoU Loss
To evaluate the alignment between predicted and ground truth bounding boxes, we adopt the Generalized Intersection over Union (GIoU) metric [54], which extends the standard IoU by incorporating shape and distance information. Given a predicted box $B_{p}$ and a ground truth box $B_{gt}$, the GIoU is defined as:
$\mathrm{GIoU}(B_{p},B_{gt})=\mathrm{IoU}(B_{p},B_{gt})-\frac{|C\setminus(B_{p}\cup B_{gt})|}{|C|}$ (3)
where
$\mathrm{IoU}(B_{p},B_{gt})=\frac{|B_{p}\cap B_{gt}|}{|B_{p}\cup B_{gt}|}$ (4)
is the standard Intersection over Union, and $C$ is the smallest enclosing box that contains both $B_{p}$ and $B_{gt}$. This formulation penalizes predictions that are far from the ground truth, even if they have partial overlap, leading to better convergence and improved spatial accuracy during training.
3. $\mathscr{L}_{\text{focal}}$: Focal Contrastive Classification Loss
To enable open-vocabulary grounding, we adopt a contrastive version of the Focal Loss [55]. Let $q_{i}$ be the $i$-th decoder query and $t_{j}$ the $j$-th token embedding in the text prompt. The model computes similarity scores via scaled dot products:
$\text{logits}_{i,j}=\frac{q_{i}^{\top}t_{j}}{\tau}$ (5)
where $\tau$ is a learned temperature parameter. For each query, we apply focal loss over the logits:
$\mathscr{L}_{\text{focal}}=-\alpha(1-p_{t})^{\gamma}\log(p_{t})$ (6)
where $p_{t}$ is the softmax probability assigned to the correct token, $\alpha=0.25$ is the class-balancing factor, and $\gamma=2.0$ is the focusing parameter (as in the original formulation).
The final loss is a weighted sum of the above components, with the following coefficients used in all experiments:
$\lambda_{\text{L1}}=0.5,\quad\lambda_{\text{GIoU}}=0.1,\quad\lambda_{\text{focal}}=0.1$
This formulation enables the model to localize and classify objects based on free-form text prompts with high precision, while preserving generalization to unseen anatomical structures and phrasing.
@@ 2.5 Implementation Details
LoRA fine-tuning was conducted using a batch size of 4 and a learning rate of $5\times 10^{-4}$ with the AdamW optimizer [58] with weight decay of $1\times 10^{-4}$. Only LoRA parameters were updated, whereas all other weights remained frozen. The training loss included focal loss for classification, L1 and GIoU for localization, and a contrastive alignment loss to optimize image-text correspondence. Early stopping was employed with a patience of 20 epochs based on validation loss to mitigate overfitting. All datasets were pooled during training, and data augmentations such as random flipping (with a probability of 50%), resizing (scale range: 0.8x to 1.2x), padding (up to 10% of the image size), erasing (with a probability of 20%, area ratio: 2–10%), and cropping (covering 80–100% of the original image) were applied. This led to a single, general-purpose model that was evaluated on all datasets (seen and unseen).
@@ 3.1 Ablation Study
Table 3 summarizes the ablation study evaluating different configurations of the Grounding DINO model combined with various segmentation heads. The baseline configuration, Grounding DINO with SAM2 (trained on natural images), shows limited segmentation performance across all datasets. Fine-tuning Grounding DINO paired with MedSAM notably improves both DSC and IoU scores. However, the best performance is achieved by fine-tuning Grounding DINO paired with SAM2, which consistently outperforms the other configurations across breast, liver, and prostate datasets. The baseline configuration (without fine-tuning) performs the worst, especially on complex datasets like AUL, where it achieves only $13.77\%$ DSC. Fine-tuning with MedSAM improves segmentation quality across all organs, but the most consistent and substantial improvements are observed when fine-tuning with SAM2. This configuration yields the best DSC and IoU scores in all datasets. For instance, on the BrEaST dataset, the DSC improves from $23.87\%$ (baseline) to $76.65\%$ when using the fine-tuned SAM2. Similarly, the MicroSeg prostate dataset shows a leap from $61.44\%$ to $88.56\%$ in DSC, with a corresponding increase in IoU to $81.17\%$. Even in the challenging AUL dataset, performance climbs to $44.02\%$ DSC, a significant gain over other methods. Statistical significance was assessed via paired t-tests across test samples. We mark results with $p<0.005$ and $p<0.01$ as ** and *, respectively. These tests confirm that the observed gains are not due to chance, particularly for breast datasets, where the differences are consistently significant. Given its consistent superiority in both mean performance and statistical robustness, we adopt fine-tuned Grounding DINO + SAM2 for all subsequent experiments.
Table 3: Ablation study comparing Grounding DINO’s segmentation performance (DSC and IoU scores) paired with different segmentation heads across multiple ultrasound datasets. The evaluated configurations include: (1) Grounding DINO + SAM2 (trained on natural images), (2) Fine-tuned Grounding DINO + MedSAM, and (3) Fine-tuned Grounding DINO + SAM2. Results are reported as mean $\pm$ standard deviation. Bolded results indicate the highest score in each comparison. The third method consistently outperforms other methods across all datasets and organs, and is used in the rest of the experiments. Statistical significance was assessed using paired t-tests: ∗∗ denotes $p<0.005$ and denotes $p<0.01$.
@@
Breast BrEaST $23.87±30^{**}$ $18.08±27^{**}$ $49.62±28^{**}$ $37.33±24^{**}$ 76.65±23 66.24±23
BUID $44.09±35^{**}$ $36.08±35^{**}$ $61.43±29^{*}$ $49.38±25^{**}$ 82.92±24 75.83±25
BUSI $29.55±34^{**}$ $23.31±31^{**}$ $51.35±30^{**}$ $39.74±26^{**}$ 77.79±26 69.29±27
Liver AUL $13.77±13^{**}$ $7.94±8^{**}$ $32.50±28$ $23.25±23^{**}$ 44.02±36 35.79±33
Prostate MicroSeg $61.44±21^{**}$ $47.47±21^{**}$ $56.84±22^{**}$ $42.75±20^{**}$ 88.56±13 81.17±15
@@ 3.2 Comparison on Seen Datasets
Table 4: Quantitative comparison of segmentation performance (DSC and IoU scores) on publicly available ultrasound datasets among SOTA methods: UniverSeg, BiomedParse, SAMUS, MedCLIP-SAM, MedCLIP-SAMv2, and the proposed method. Results are reported as mean $\pm$ standard deviation. Bolded results indicate the highest score in each comparison. Statistical significance was assessed using paired t-tests: ∗∗ denotes $p<0.005$ and denotes $p<0.01$.
@@
Breast BrEaST $60.30±23$ $47.01±24^{**}$ $76.36±20$ $65.19±22$ $60.78±22$ $46.95±21$ $65.02±21$ $51.18±19^{**}$ $72.91±25$ $62.24±25$ 76.65±23 66.24±23
BUID $56.30±30^{**}$ $45.27±29^{**}$ $82.77±21$ $74.54±23$ $71.74±17$ $57.93±16$ $70.13±22^{*}$ $57.90±24^{**}$ $78.48±23$ $69.48±26$ 82.92±24 75.83±25
BUSUC $51.94±28^{**}$ $39.89±26^{**}$ $87.17±12^{**}$ $78.92±16^{**}$ $71.46±17^{**}$ $58.06±19^{**}$ $80.66±11^{**}$ $68.88±14^{**}$ $87.40±11^{**}$ $78.99±15^{**}$ 91.68±6 85.16±9
BUSUCLM $52.96±22^{**}$ $39.19±21^{**}$ $66.34±31$ $56.50±30$ $44.34±32^{**}$ $34.32±29^{**}$ $56.88±21^{**}$ $42.48±19^{**}$ $66.09±27$ $54.66±27$ 76.74±24 66.75±24
BUSB $82.58±15$ $72.87±20$ $73.93±31$ $66.12±31$ $64.40±20^{**}$ $50.52±20^{**}$ $73.84±23^{**}$ $62.84±24^{**}$ $82.99±23$ $75.46±23$ 91.04±8 84.30±11
BUSI $43.19±26^{**}$ $31.16±22^{**}$ 89.39±14** 82.72±15** $82.45±10^{**}$ $71.19±13^{**}$ $56.64±30^{**}$ $45.51±29^{**}$ $63.60±34^{**}$ $54.40±32^{**}$ $77.79±26$ $69.29±27$
STU $64.73±17$ $50.16±19$ $87.07±5$ $77.50±8^{**}$ $68.46±11$ $53.10±12$ $83.77±9$ $73.03±13$ 90.11±6** 82.56±10** $87.89±6$ $78.87±9$
S1 $57.30±25^{**}$ $44.25±24^{**}$ $93.28±4$ $87.63±6$ $76.09±20$ $64.54±20$ $78.92±9^{**}$ $66.02±12^{**}$ $88.04±8^{**}$ $79.40±11^{**}$ 93.37±5** 87.96±8**
Thyroid TN3K $39.46±21^{**}$ $26.88±18^{**}$ $56.96±34^{**}$ $47.07±31^{**}$ 80.96±12** 69.43±14** $44.42±27^{**}$ $32.60±23^{**}$ $43.28±31^{**}$ $32.92±27^{**}$ $72.20±26$ $61.77±26$
TG3K $37.27±19^{**}$ $24.57±14^{**}$ $1.57±5^{**}$ $0.88±3^{**}$ 82.95±6** 71.22±8** $67.82±17^{**}$ $53.50±17^{**}$ $70.75±20^{**}$ $57.73±20^{**}$ $80.32±9$ $68.04±12$
@@
AUL $14.52±15^{**}$ $8.60±10^{**}$ $19.78±32^{**}$ $15.48±26^{**}$ $28.18±22$ $18.59±17$ $32.12±27$ $22.69±22^{**}$ $32.86±34$ $25.48±29$ 44.02±36 35.79±33
Prostate MicroSeg $11.62±9^{**}$ $6.44±6^{**}$ $23.98±24^{**}$ $16.25±20^{**}$ $70.84±18^{**}$ $57.28±18^{**}$ $58.67±15^{**}$ $42.94±14^{**}$ $53.04±30^{**}$ $41.10±25^{**}$ 88.56±13 81.17±15
RegPro $16.84±14^{**}$ $9.89±9^{**}$ $38.88±32^{**}$ $29.53±27^{**}$ $49.22±29^{**}$ $37.30±25^{**}$ $42.90±28^{**}$ $31.26±23^{**}$ $46.90±31^{**}$ $36.00±27^{**}$ 70.77±31 61.63±29
Kidney KidneyUS $22.22±15^{**}$ $13.36±10^{**}$ $48.03±23^{**}$ $34.55±20^{**}$ $56.37±18^{**}$ $41.39±18^{**}$ $59.86±16^{**}$ $44.45±15^{**}$ $64.93±21^{**}$ $51.10±20^{**}$ 73.49±16 60.48±19